Bailey Corporation is considering modernizing its production by purchasing a new machine and selling an old machine. The following data have been collected on this investment:



The income tax rate is 40%, and the required rate of return is 16%. Amortization is $5,000 per year for the old machine. The new machine would be amortized $7,600 in 20x1, $5,700 in 20x2, $3,800 in 20x3, and $1,900 in 20x4. Assume Bailey would purchase the new machine in December 20x0 and dispose of the old machine in January 20x1.

The tax effect of selling the new machine in 20x4 would be:
Answer #1

Calculation of tax effect of selling new machine in 2004

A. Increase in tax amount due to decrease in depreciation expenses in the year 2004

=(Depreciation on old machine - depreciation on new asset) x Tax rate

=(5000-1900) x 40%

=1240

​​B. Increase in Tax on sale proceeds of machinery

=(Salvage value of new machinery-salvage value of old machinery) x 40%

=(5000-0)

=2000

C.​​​​​Increase in tax due to decrease in operating costs

=(Operating cost for old machine- operation costs for new machine) x 40%

=(18000-14000) x 40%

=1600.

So total increase in tax in year 2004 due to buying of new machine is

=1240+2000+1600

=$4840